OpenStudy (anonymous):

solve this system by graphing, is it consistant,inconsistant, dependant or independant 8x-7y=35 7y-8x=-35

OpenStudy (saifoo.khan):

consistent and dependent.

OpenStudy (saifoo.khan):

Infinite solutions.

jimthompson5910 (jim_thompson5910):

8x-7y=35 7y-8x=-35 8x-7y=35 -8x+7y=-35 Now add the equations to get (8x-7y)+(-8x+7y) = 35+(-35) (8x+(-8x)) + (-7y+7y) = 0 0x+0y = 0 0 = 0 Since the equation is ALWAYS true, this means that there are an infinite number of solutions. So the system is consistent and dependent.

OpenStudy (anonymous):

Beam me up Jim ,I'm on your ship and medals to both you and saifoo!